# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 2. Preliminaries
$w_{i}\cdot g=(w_{i}\cdot h_{i}\cdot w_{i}^{-1})\cdot u\cdot p_{1}\cdot p_{2}\cdot p_{3}\cdot x\cdot\gamma.$
Since $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ and $R_{u}(\mathbf{P}_{0})$ fix every element in $V_{\beta_{0}}$, and $\mathbf{T}$ and $\mathbf{M}_{a}$ preserve the weight space $V_{\beta_{0}}$, we obtain
$\rho(w_{i}\cdot g)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2}\cdot p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2})(\rho(p_{3}\cdot x)\mathbb{Z}^{d}\cap V_{\beta_{0}}).$
The lemma then follows from the facts that $p_{1}\in\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$, $p_{2}$ is in a fixed compact subset in $\mathbf{M}_{a}(\mathbb{R})$ and $\rho(p_{3}\cdot x)\mathbb{Z}^{d}$ is commensurable with $\mathbb{Z}^{d}$ by $\Gamma$ and $\mathcal{K}$. ∎
Now we consider the case that $w_{i}=e$ is the identity element in ${}_{\mathbb{Q}}W$. Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $\mathbf{F}_{w_{i}}=\mathbf{F}_{e}$ and $\mathbf{H}_{w_{i}}=\mathbf{H}_{e}$. By Lemma 2.4, there exist $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that
@@
where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$. Then we can compute the denominator $d(g)$ of $g$ as follows:
@@
$d(g)=c_{x}\cdot e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}$
@@
for some $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x_{j}\in\mathcal{K}$ and $\gamma\in\Gamma$.
3. Counting rational points
In this section, we consider the problem of counting rational points in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. To estimate the Hausdorff dimension, one usually constructs boxes centered at certain rational points and obtains a tree-like subset. At each level of the tree-like subset, the boxes are disjoint. In our case, it may happen that the boxes at each level are not disjoint if we proceed in the usual way. It implies that the boxes constructed are too many, and we have to sieve out some proportion of the rational points so that the boxes centered at the remaining rational points are disjoint. It is also required that the number of the remaining rational points is not small compared to the total number of rational points in order to avoid any loss of Hausdorff dimension. For our purpose, in the following, we will first sieve some proportion of rational points according to several algebraic conditions, and then count the remaining rational points by the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$. The disjointness property of the boxes centered at the remaining rational points will follow from the transversal structure of some submanifolds in the homogeneous space $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
@@ 3. Counting rational points
The connected component of identity of $\mathbf{L}$ in the Zariski topology is denoted by $\mathbf{L}^{0}$, and the connected component of identity in the Lie group $\mathbf{L}(\mathbb{R})$ is denoted by $\mathbf{L}(\mathbb{R})^{0}$. We write $A\lesssim B$ $(A\gtrsim B)$ if there exists a constant $c>0$ such that
$A\leq c\cdot B\quad(A\geq c\cdot B).$
If $A\lesssim B$ and $A\gtrsim B$, then we write $A\sim B$. We will specify the implicit constants in the contexts if necessary.
In this section and next section, we assume that the action of the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. This holds true when the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into any $\mathbb{Q}$-simple factor of $\mathbf{G}$. Later in §5, we will explain how to establish the results of counting rational points and Hausdorff dimension estimates in the case that $\{a_{t}\}_{t\in\mathbb{R}}$ appears only in some $\mathbb{Q}$-simple factors of $\mathbf{G}$.
Let $U$ be an open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. For any $0<A<B$, define
$S(U,A,B)=\{q\in U:q\textup{ rational and }A\leq d(q)\leq B\}.$
We know that the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ has the following decomposition
@@
$\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0.$
Denote by $\pi_{\ker(\beta_{0})}$ the projection from $\mathfrak{a}$ to $\ker(\beta_{0})$ according to this decomposition. We also write $\pi_{\ker(\beta_{0})}(a)$ for $\pi_{\ker(\beta_{0})}(\log(a))$ whenever $a\in\exp(\mathfrak{a})$. Here $\log$ is the inverse of the exponential map $\exp$. For any compact subset $L$ in $\operatorname{Lie}(a_{t})$ and any compact subset $K$ in $\ker(\beta_{0})$, we denote by
@@
For any $x_{j}\in\mathcal{K}$, any compact subset $K_{1}$ in $\ker(\beta_{0})$, any compact subset $K_{2}$ in $\mathbf{H}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$, we define
@@
to be the set of all rational points $q$ in $U$ such that
@@
then $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$ (and such elements exist as $e\in\mathcal{K}$). The elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ are the rational points we are interested in when we construct a tree-like subset in §4. We denote by
@@
To count the rational points in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$, we need results about limiting distributions of translates of unipotent orbits pushed by the one-parameter semisimple flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. The following is a direct consequence of the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
@@
Let $x\in\mathbf{G}(\mathbb{R})^{0}/\Gamma$ and $W\subset\mathbf{G}(\mathbb{R})^{0}/\Gamma$ an open bounded subset whose boundary has measure zero with respect to the invariant probability measure $\mu_{\mathbf{G}^{0}(\mathbb{R})/\Gamma}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Let $\chi_{W}$ denote the characteristic function of $W$. Then for any bounded open subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})$ we have
[…]
@@
Note that $R_{u}(\overline{\mathbf{P}}_{0})$ is not necessarily the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$. Here we can still apply the mixing property to obtain Proposition 3.1 as long as $R_{u}(\overline{\mathbf{P}}_{0})$ is contained in the group generated by the unstable subgroup and the central subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$. (See [25, §2].)
Now we can follow the same arguments as in [17, 48] and measure the size of the subset $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ for a fixed index $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$, and for some sufficiently large numbers $B>A>0$. Recall that
@@
where $\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\cap\mathbf{P}_{\beta_{0}}$. We fix a Haar measure $\mu_{\mathbf{H}_{e}}$ on $\mathbf{H}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\mathbf{F}_{e}}$ on $\mathbf{F}_{e}(\mathbb{R})$ such that the product maps
$\mathbf{F}_{e}\times\mathbf{H}_{e}\to\mathbf{F}_{e}\cdot\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\textup{ and }\mathbf{H}_{e}\times\mathbf{F}_{e}\to\mathbf{H}_{e}\cdot\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{0})$
induce Haar measures on $R_{u}(\overline{\mathbf{P}}_{0}).$ Then for any $q\in\mathbf{F}_{e}(\mathbb{Q})$, we define $m_{\mathbf{H}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which is supported on $\mathbf{H}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\mathbf{H}_{e}}$ via the product map
$\mathbf{H}_{e}\times\{q\}\to\mathbf{H}_{e}\cdot q\subset R_{u}(\overline{\mathbf{P}}_{0}).$
We define
@@
Note that $m_{\mathbf{H}_{e}}$ is not a locally finite measure on $R_{u}(\overline{\mathbf{P}}_{0})$, and it is defined by the leaf-wise measures on the countable leaves through the rational elements in $\mathbf{F}_{e}(\mathbb{Q})$ in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ induced by the group action of $\mathbf{H}_{e}$ on $R_{u}(\overline{\mathbf{P}}_{0})$. Note also that by Corollary 2.3, we have
@@
In the following, we estimate the size of the subset
@@
with respect to the measure $m_{\mathbf{H}_{e}}$ for sufficiently large $l>0$. For convenience, we write
@@
Let $q$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, we may write
@@
for some $a\in\mathbf{T}(\mathbb{R})$, $h\in\mathbf{H}_{e}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$, $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, $x_{k}\in C$ and $\gamma\in\Gamma$. Then
@@
$\iff$ $q\in U,\;A_{l}\leq d(q)\leq B_{l},\textrm{ and }a\in\exp(\mathfrak{a}),\pi_{\ker(\beta_{0})}(a)\in K_{1},h\in K_{2},m\in K_{3},u\in K_{4},k=j.$
@@
Since $\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma$ is a compact subset in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$, there exist $\delta_{0}>0$ and a small neighborhood of identity $B_{\mathbf{F}_{e}}(\delta_{0})$ in $\mathbf{F}_{e}(\mathbb{R})$ such that
@@
is a homeomorphism. We conclude that the following product map
$B_{\mathbf{F}_{e}}(\delta_{0})\times a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$
gives a transversal structure near the set $a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma$ in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Consequently, for any $p,q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, the subsets $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot p\Gamma$ and $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot q\Gamma$ are disjoint, where
@@
Now we estimate the size of $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ with respect to the measure $m_{\mathbf{H}_{e}}$. First we prove an upper bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$. Fix a sufficiently small number $0<\epsilon<\delta_{0}$ such that
@@
Since $B_{\mathbf{F}_{e}}(\delta_{0},T)$ and $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$ form a transversal structure near $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, by Proposition 3.1, we deduce that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))$
[…]
$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\cdot\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}(B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{K_{1},I_{0}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma)$
$\sim$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))$
@@
where the implicit constant depends only on $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $I_{0}$, $\mathbf{G}$ and $\Gamma$. Similarly, using the arguments in [17, §4], we can prove a lower bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$ and for any sufficiently large $l>0$ we have

### v2

@@ 2. Preliminaries
$w\cdot g=(w\cdot h\cdot w^{-1})\cdot u\cdot p_{1}\cdot p_{2}\cdot p_{3}\cdot x\cdot\gamma.$
Since $w\mathbf{H}_{w}w^{-1}$ and $R_{u}(\mathbf{P}_{0})$ fix every element in $V_{\beta_{0}}$, and $\mathbf{T}$ and $\mathbf{M}_{a}$ preserve the weight space $V_{\beta_{0}}$, we obtain
$\rho(w\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2}\cdot p_{3}\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}=\rho(p_{1}\cdot p_{2})(\rho(p_{3}\cdot x)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}).$
The lemma then follows from the facts that $p_{1}\in\mathbf{T}$ acts as scalars in $V_{\beta_{0}}$, $p_{2}$ is in a fixed compact subset in $\mathbf{M}_{a}(\mathbb{R})$ and $\rho(p_{3}\cdot x)\cdot\mathbb{Z}^{d}$ is commensurable with $\mathbb{Z}^{d}$ by $\Gamma$ and $\mathcal{K}$. ∎
Now we consider the case where $w=e$ is the identity element in $_{\mathbb{Q}}\mathcal{W}$. Then $\mathbf{F}_{w}=\mathbf{F}_{e}$ and $\mathbf{H}_{w}=\mathbf{H}_{e}$. Let $g$ be a rational element in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, there exist $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that
@@
where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$. Then we can compute the height $0pt(g)$ of $g$ as follows:
@@
$0pt(g)=c_{x}\cdot e^{\beta_{0}(a)\cdot\dim V_{\beta_{0}}}$
@@
for some $h\in\mathbf{H}_{e}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x_{j}\in\mathcal{K}$ and $\gamma\in\Gamma$. Furthermore, if $p=a\cdot m\cdot u$ where $a\in\mathbf{T}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$ and $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, then $a$ is called the central coordinate of $g$ which we denote by $a_{g}$.
3. Counting rational elements
In this section, we consider the problem of counting rational elements in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. As discussed in §1, to get a lower bound of the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, we need to select certain rational elements and construct a Cantor type subset from disjoint neighborhoods of these rational elements. For our purpose, in the following, we will select rational elements according to several algebraic conditions, and then count these rational elements by the mixing property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. The disjointness of the neighborhoods around these rational elements will follow from the transversal structure of some submanifolds in $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
@@ 3. Counting rational elements
The identity component of $\mathbf{L}$ with respect to the Zariski topology is denoted by $\mathbf{L}^{0}$, and the identity component of $\mathbf{L}(\mathbb{R})$ (as a Lie group) is denoted by $\mathbf{L}(\mathbb{R})^{0}$.
In this section and §4, we assume that the action of the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. This assumption holds when $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into any $\mathbb{Q}$-simple factor of $\mathbf{G}$. Later in §5, we will explain how to establish the results about counting rational elements and estimating Hausdorff dimensions in the case where $\{a_{t}\}_{t\in\mathbb{R}}$ projects trivially into some of the $\mathbb{Q}$-simple factors of $\mathbf{G}$.
Let $U$ be a small open bounded subset in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which projects injectively into $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$. For any $0<A<B$, define
$S(U,A,B)=\{q\in U:q\textup{ rational and }A\leq 0pt(q)\leq B\}.$
The Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ has the following direct sum decomposition
@@
$\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0\;(t>0).$
Denote by $\pi_{\ker(\beta_{0})}$ the projection of $\mathfrak{a}$ onto $\ker(\beta_{0})$ along the linear subspace $\operatorname{Lie}(a_{t})$. For convenience, we will write $\pi_{\ker(\beta_{0})}(a)$ for $\pi_{\ker(\beta_{0})}(\log(a))$ whenever $a\in\exp(\mathfrak{a})$ (here $\log$ is the inverse of the exponential map $\exp$). For any compact subset $L$ in $\operatorname{Lie}(a_{t})$ and any compact subset $K$ in $\ker(\beta_{0})$, we denote by
@@
For any $x_{j}\in\mathcal{K}$ ($j\in J$), any compact subset $K_{1}$ in $\ker(\beta_{0})$, any compact subset $K_{2}$ in $\mathbf{H}_{e}(\mathbb{R})$, any compact subset $K_{3}\subset\mathbf{M}_{a}(\mathbb{R})\cap\mathbf{G}(\mathbb{R})^{0}$ and any compact subset $K_{4}\subset R_{u}(\mathbf{P}_{0})(\mathbb{R})$, we define
@@
to be the set of all rational elements $q$ in $U$ such that
@@
then $x_{j}\in\mathbf{G}(\mathbb{R})^{0}$ (and such elements exist as $e\in\mathcal{K}$). The elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ are the rational elements we select when we construct a Cantor type subset in §4. We denote by
@@
To count the rational elements in $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ $(j\in J)$, we need the following result about limiting distributions of translates of unipotent orbits pushed by the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$, which is a direct consequence of the mixing property of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$.
@@
Let $x\in\mathbf{G}(\mathbb{R})^{0}/\Gamma$ and $W\subset\mathbf{G}(\mathbb{R})^{0}/\Gamma$ an open bounded subset whose boundary has measure zero with respect to the invariant probability measure $\mu_{\mathbf{G}^{0}(\mathbb{R})/\Gamma}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$. Let $\chi_{W}$ denote the characteristic function of $W$. Then for any bounded open subset $U$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ we have
$\lim_{t\to\infty}\frac{1}{\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)}\int_{U}\chi_{W}(a_{t}\cdot nx)d\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(n)=\int_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}\chi_{W}d\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}$
@@
Note that $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is not necessarily the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}(\mathbb{R})$. Here we can still apply the mixing property to obtain Proposition 3.1 as long as $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ is contained in the group generated by the unstable subgroup and the centralizer of $\{a_{t}\}_{t\in\mathbb{R}}$ in $\mathbf{G}(\mathbb{R})$. (See [49, §2].)
Now we define a measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, which will be used to study the size of the subset $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A,B)$ for some sufficiently large numbers $0<A<B$. Recall that
@@
where $\mathbf{F}_{e}=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $\mathbf{H}_{e}=R_{u}(\overline{\mathbf{P}}_{0})\cap\mathbf{P}_{\beta_{0}}$. By [55, Proposition 5.26], we fix a Haar measure $\mu_{\mathbf{H}_{e}}$ on $\mathbf{H}_{e}(\mathbb{R})$ and a Haar measure $\mu_{\mathbf{F}_{e}}$ on $\mathbf{F}_{e}(\mathbb{R})$ such that the product maps
[…]
induce Haar measures on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ For any $q\in\mathbf{F}_{e}(\mathbb{Q})$, we define $m_{\mathbf{H}_{e}q}$ to be the locally finite measure defined on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which is supported on $\mathbf{H}_{e}(\mathbb{R})\cdot q$ and induced by $\mu_{\mathbf{H}_{e}}$ via the product map
$\mathbf{H}_{e}(\mathbb{R})\times\{q\}\to\mathbf{H}_{e}(\mathbb{R})\cdot q\subset R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$
Then we define
@@
Note that $m_{\mathbf{H}_{e}}$ is not a locally finite measure on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, and it is defined by the leaf-wise measures on the countable leaves $\mathbf{H}_{e}(\mathbb{R})\cdot q$ $(q\in\mathbf{F}_{e}(\mathbb{Q}))$ in the foliation $\mathcal{F}_{\mathbf{H}_{e}}$ induced by the group action of $\mathbf{H}_{e}(\mathbb{R})$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Note also that by Corollary 2.3, we have
@@
In the following, we use the arguments in [25, 82] and estimate the $m_{\mathbf{H}_{e}}$-measure of the subset
@@
for $j\in J$ with $x_{j}\in\mathbf{G}(\mathbb{R})^{0}\cap\mathcal{K}$ and for sufficiently large $l>0$. For convenience, we write
@@
Let $q$ be a rational element in $U\subset R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Lemma 2.4, we may write
@@
for some $a\in\mathbf{T}(\mathbb{R})$, $h\in\mathbf{H}_{e}(\mathbb{R})$, $m\in\mathbf{M}_{a}(\mathbb{R})$, $u\in R_{u}(\mathbf{P}_{0})(\mathbb{R})$, $x_{k}\in\mathcal{K}$ and $\gamma\in\Gamma$. Then
@@
$\iff$ $q\textup{ rational in }U,\;A_{l}\leq 0pt(q)\leq B_{l},\textrm{ and}$
$\quad a\in\exp(\mathfrak{a}),\pi_{\ker(\beta_{0})}(a)\in K_{1},h\in K_{2},m\in K_{3},u\in K_{4},k=j.$
@@
$a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma=a_{T}\cdot U\Gamma/\Gamma\cap\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma.$
Since $\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}$ is a compact subset in $\mathbf{P}_{\beta_{0}}(\mathbb{R})$ and
$\mathbf{F}_{e}(\mathbb{R})\cap\mathbf{P}_{\beta_{0}}(\mathbb{R})=R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})\cap\mathbf{P}_{\beta_{0}}(\mathbb{R})=\{e\},$
there exists a small neighborhood of identity $B_{\mathbf{F}_{e}}(\delta_{0})$ in $\mathbf{F}_{e}(\mathbb{R})$ (for some $\delta_{0}>0$) such that
@@
is a homeomorphism, and hence the following product map
(1) $B_{\mathbf{F}_{e}}(\delta_{0})\times a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0})\cdot a_{T}\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma$
is bijective. Consequently, for any $p,q\in S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, the subsets $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot p\Gamma$ and $B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot q\Gamma$ are disjoint, where
@@
Now we estimate the $m_{\mathbf{H}_{e}}$-measure of $S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})$, following the arguments in [25, §4]. First we prove an upper bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$. Fix a sufficiently small number $0<\epsilon<\delta_{0}$ such that
@@
Since the map (1) is bijective, the following product map
$B_{\mathbf{F}_{e}}(\delta_{0},T)\times S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma\to B_{\mathbf{F}_{e}}(\delta_{0},T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l})\Gamma/\Gamma$
is also bijective. By Proposition 3.1 and the fact that $U$ projects injectively into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap\Gamma$, we deduce that
$m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon,T))$
$=$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon,T)\cdot S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$
$\leq$ $\int_{B_{\mathbf{F}_{e}}(\epsilon)\cdot U}\chi_{B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma}(a_{T}u\Gamma)d\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(u)$
$\asymp$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(B_{\mathbf{F}_{e}}(\epsilon)\cdot U)\cdot\mu_{\mathbf{G}(\mathbb{R})^{0}/\Gamma}(B_{\mathbf{F}_{e}}(\epsilon)\exp(\mathfrak{a}_{I_{0},K_{1}})K_{2}K_{3}K_{4}x_{j}\Gamma/\Gamma)$
$\asymp$ $\mu_{R_{u}(\overline{\mathbf{P}}_{0})}(U)\cdot\mu_{\mathbf{F}_{e}}(B_{\mathbf{F}_{e}}(\epsilon))$
@@
where the implicit constant depends only on $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $I_{0}$, $\mathbf{G}$ and $\Gamma$. Similarly, using the arguments in [25, §4], we can prove a lower bound for $m_{\mathbf{H}_{e}}(S_{K_{1},K_{2},K_{3},K_{4},j}(U,A_{l},B_{l}))$ and for any sufficiently large $l>0$ we have